Tom is 3 years older than sara . Three years ago sara's age was four fifth tom's age . Find their present age.
Answer by Alan3354(69443)

Tom is 3 years older than sara . Three years ago sara's age was four fifth tom's age . Find their present age.
----------
3 yrs ago:
S = 4T/5
S = T - 3
---
T-3 = 4T/5
5T-15 = 4T
T = 15
==========
Tom is 18
Sara is 15